### v1

# IDs for AI Systems

Abstract
AI systems are increasingly pervasive, yet information needed to decide whether and how to engage with them may not exist or be accessible. A user may not be able to verify whether a system satisfies certain safety standards. An investigator may not know whom to investigate when a system causes an incident. A platform may find it difficult to penalize repeated negative interactions with the same system. Across a number of domains, IDs address analogous problems by identifying particular entities (e.g., a particular Boeing 747) and providing information about other entities of the same class (e.g., some or all Boeing 747s). We propose a framework in which IDs are ascribed to instances of AI systems (e.g., a particular chat session with Claude 3), and associated information is accessible to parties seeking to interact with that system. We characterize IDs for AI systems, argue that there could be significant demand for IDs from key actors, analyze how those actors could incentivize ID adoption, explore potential implementations of our framework, and highlight limitations and risks. IDs seem most warranted in high-stakes settings, where certain actors (e.g., those that enable AI systems to make financial transactions) could experiment with incentives for ID use. Deployers of AI systems could experiment with developing ID implementations. With further study, IDs could help to manage a world where AI systems pervade society.
1 Introduction
AI systems are becoming increasingly commonplace. While current systems can struggle to complete complex tasks (Mialon et al. 2023; Liu et al. 2023; Kinniment et al. 2023; Xie et al. 2024; Jimenez et al. 2024), capabilities seem likely to improve (Hoffmann et al. 2022; Epoch 2023; Erdil and Besiroglu 2023; Ho et al. 2024), and future AI agents could carry out a broad range of tasks with only minimal human intervention (Chan et al. 2023; Shavit et al. 2023; Chan et al. 2024). Several commercially deployed AI systems can already search the web, send emails, and write code (OpenAI 2024b; Anthropic 2024a). Even when they do not function reliably, AI systems might still be widely used, whether because of cost advantages, hype, or the externalization of their harms (De La Garza 2020; BBC 2021; Raji et al. 2022).
At the same time, information to make decisions about engaging with AI systems may not exist or be accessible. Although there may be obligations to inform parties that they are interacting with an AI system (e.g., EU AI Act Article 50.1 (Parliament 2024)), those parties may not know with which AI system they are interacting. For example, a party that knowingly interacts with an AI system may not be aware that their system is relatively more vulnerable to adversarial attacks (Zhan et al. 2024). Awareness of such elevated risk could justify additional precautions, such as reviewing the system’s actions or abstaining from interaction altogether. Furthermore, the same party—or an investigator—may lack the information to pursue recourse if the system causes harm. Information about the system or the interaction (e.g., whether a system behaved according to safety standards, or the identity of the deployer) could aid incident investigation, allocation of liability, or other legal action (Buiten, de Streel, and Peitz 2023; Buiten 2024; Kolt 2024; Wills 2024).
Across numerous domains, IDs held by software, assets, individuals, and organizations address analogous problems. IDs can help to ascertain compliance with standards or regulation. For instance, an individual aircraft’s tail number is associated to the aircraft’s incident and maintenance history, which could inform safety assessments from regulators or aircraft operators. Furthermore, IDs can help to establish whether trust is warranted. A website’s valid HTTPS certificate assures users of the website domain’s authenticity, and provides a way for users to establish a secure communication channel. [^1] Despite a lack of centralized enforcement, HTTPS gradually became the norm due to widespread awareness of its security benefits and collective advocacy from web browsers, search engines, and other organizations (Encrypt 2024; Hancock 2021). Finally, IDs can facilitate redress. Serial numbers on consumer products enable customer support, product recalls, and attribution of liability. A key feature that enables all of the above functions is that IDs are specific to particular entities (e.g., a particular Boeing 747), although they may also contain information about other entities of the same class (e.g., some or all Boeing 747s).
Identifying particular AI systems, which we term instances, could be similarly useful. An instance corresponds to a context window and a user. [^2] This definition of instance does not take memory into account and is anchored upon current language models. While it will suffice for this work, see Appendix A for a more general definition. For example, a particular user’s chat session with ChatGPT (with e.g. a GPT-4 backend) is an instance, and that instance could interact with other parties by sending an email through the user’s account (OpenAI 2024b). Separate instances can behave differently, whether because of user instructions (Shanahan, McDonell, and Reynolds 2023; Wei et al. 2023; Bai et al. 2022; Zou et al. 2023; Agarwal et al. 2024), hijacking by an attacker (Greshake et al. 2023; Zhan et al. 2024), or malfunction. As such, instance-specific information could aid decisions about interactions. For example, repeated negative interactions with the same instance may cause a party to reject further interaction from the instance. The ability to distinguish instances from one another could aid incident investigation (e.g., a client files a complaint about an instance) and allocation of liability (e.g., if a third party hijacked the instance). Moreover, instance-level identification would allow flexible grouping of instances. For example, an incident investigator could be interested in all the instances that interacted with a particular party.
Yet, current mechanisms do not identify instances. System documentation (Gebru et al. 2021; Mitchell et al. 2019; Gilbert et al. 2023; Bommasani et al. 2023) provides information about systems, but such documentation is not an ID for instances or systems (however, an ID could include such documentation). API tokens for services, such as for hotel booking (OpenAI 2023a), do identify entities that use the tokens, but are usually only user- or device-specific. Even if API requests do include a string with the name of the system, any actor could mimic the string in response to the same service. Finally, user accounts only separate the activities of AI systems from different users.
To prepare for a world with ubiquitous AI interactions, we propose (instance-level) IDs for AI systems. An AI ID is a container for 1) an identifier and 2) attributes. An identifier is a unique string that refers to an instance. [^3] It may be desirable in the future to have identifiers that refer to even more granular parts of an AI system. For example, one could have an identifier refer to an instance as it operated between times $t$ and $t+10$. An identifier could be randomly generated, or could itself encode some information about the instance (e.g., the identifier could include the exact time at which the instance started operating). An attribute is any information that could pertain to the instance, and could include behaviour, properties, context, or relationships to other instances or systems. An identifier enables the association of attributes to an instance, similar to how serial numbers associate information to a particular product. Attributes could be specific to an instance (e.g., prior incidents associated with the instance), or could apply more broadly to other instances (e.g., a system [^4] A system (e.g., ChatGPT) could use different models (e.g., GPT-4, GPT-3.5) as a backend. We consider system cards to include model cards (Mitchell et al. 2019). card). An ID could directly include attributes (e.g., the name of the deployer) or could link to them (e.g., a link to a database of prior incidents).
Different instances or systems could warrant attributes of differing levels of granularity or detail. For example, attaching a user identifier to an ID may only be appropriate in high-stakes settings, such as when a company uses an AI system to interface with critical infrastructure. Prior incidents associated with an instance would likely be more useful for future AI agents (Chan et al. 2023; Shavit et al. 2023; Chan et al. 2024; Gabriel et al. 2024), which could act autonomously and persist over long durations.
Since IDs would contain information about AI systems, they could complement measures that attempt to verify whether AI systems—or their artifacts—are present. Watermarks (Liu et al. 2024; Wang et al. 2021) embed origin information in AI outputs, while content provenance measures (C2PA 2023) embed such information in metadata. Both types of techniques could embed IDs. Other measures, like CAPTCHAs (Shet 2014), verify that a human is performing an action, so as to reduce service abuse. If AI systems become essential for many tasks, such as web search and account registration, it may be useful to allow them to bypass CAPTCHAs in exchange for presenting an ID. [^5] Advances in AI capabilities may render CAPTCHAs ineffective. We discuss this possibility further in Section 4. Parties could use this ID to track and disincentivize abuse.
1.1 Contributions
We propose IDs for AI systems. First, we characterize the central properties of IDs. Second, we argue that there will likely be demand for IDs from several key actors, especially in high-stakes settings. We also explore potential ways for these actors to incentivize ID use. Third, we analyze potential technical implementations of IDs. Finally, we investigate some of the limitations of IDs and of our analysis, including privacy and security risks of IDs, as well as uncertainty about the broader societal consequences of IDs for AI systems.
We recommend limited experimentation with IDs. IDs seem most warranted in high-stakes settings, where certain actors could experiment with incentives for ID use. For example, actors that provide interfaces for AI systems to carry out financial transactions could impose rate limiting whenever IDs are not present. However, instances without IDs should still be allowed to access services. Deployers of AI systems could experiment with developing the technical infrastructure to enable IDs. ID implementation in decentralized deployments will require more study.
Figure 1: IDs contain a unique identifier along with attributes (e.g., a system card or [a link to] previous incidents). We also display some potential actions that parties might take based on information in an ID.
2 Definitions
We collate additional definitions that will be useful for the rest of this work.
An instance of an AI system is an abstraction that corresponds to a context window and an (initial) user. Different instances are causally independent from each other, unless they interact or affect a shared entity (e.g., the same user’s bank account) in the world. Except where we point out additional nuances, readers can consider an instance to be roughly the same as a chat session with a chatbot. An ID for an instance would thus identify that session to other parties interacting with the chatbot, actors investigating an incident that the chatbot caused, etc. We provide here some examples of instances in the context of chat sessions, and defer a more general definition and discussion of its limitations to Appendix A:
  • A user’s continuous chat session, without regenerating responses, is an instance.
  • In the context of a chat session, regenerating a response creates a new instance.
A party is any entity that interacts with, or is deciding whether to interact with, an AI system. Categories of parties include humans, organizations, or computer programs.
A deployer is an organization that runs AI systems for users. As of April 2024, examples of deployers include Microsoft, OpenAI, Anthropic, and Cohere. Developers can be, but are not necessarily, deployers. For example, OpenAI, Anthropic, Google, and Cohere develop and deploy their own systems. Although chains of deployers can exist (e.g., deployer $A$ runs a system for deployer $B$, who modifies that system and serves it to user $C$), we leave deeper consideration of this nuance for future work. When we refer to a centralized deployment setting, we mean a setting where deployers run AI systems. When we refer to a decentralized deployment setting, we mean a setting where users run AI systems for themselves, whether on their own or rented (e.g., cloud compute) hardware.
A service is software that allows an entity to perform tasks. For example, software which an individual uses to perform online banking is a service. An (AI) plugin (OpenAI 2023a; Richards 2023) is software that allows AI systems to interact with services. For instance, plug-ins allow GPT-4 to interact with web search, Wikipedia, and Twitter (Significant-Gravitas 2024). A plugin developer is not necessarily the same as a service provider—the actor that develops and maintains the service—since the former can take advantage of existing interfaces for services. Third parties (plugin developer) use Microsoft’s (service provider) existing software interface for Bing to allow GPT-4 to perform web searches (Richards 2023).
3 Characterizing IDs
An ID is a container (Korenhof et al. 2014) for 1) an identifier (that corresponds to an instance) and 2) attributes. We characterize a design space for IDs: the attributes it includes, to whom it is accessible, and to what extent it is verifiable. Our goal in this section is to investigate possible designs and why they may be desirable, rather than to prescribe specific choices.
3.1 Attributes
In addition to the identifier, an ID can contain (or link to) attributes: any information that could be useful to a party interacting with the corresponding instance. An attribute can vary along two dimensions.
Category: Categories of attributes include (but are not limited to) behaviour, properties, context, and relationships to other instances or systems. Some examples are:
  • Behaviour: prior incidents (Wei and Heim 2024)
  • Properties: information in a system card (Mitchell et al. 2019; Gebru et al. 2021); the results of evaluations (Shevlane et al. 2023; Weidinger et al. 2023)
  • Context: the system prompt; external memory (Wang et al. 2023)
  • Relationships to other instances or systems: any other AI systems the instance has created or is running; any instances that have created the instance in question
Behaviour and properties could straightforwardly inform AI interaction decisions. Context could inform parties about potential behaviour (e.g., there is a jailbreak in the system prompt or external memory (Cohen, Bitton, and Nassi 2024)). Lastly, relationship information could aid incident investigation and response. Suppose a user instructs an instance to carry out a personalized influence campaign. The instance could create many descendant instances, each of which could target an individual person. If the ID of a descendant instance is linked to the ID of the original, ancestor instance, it could be easier to investigate and resolve such misuse. In Sections 5 and 6, we discuss when this ID linking is possible.
Specificity: An attribute could provide information about multiple instances (even of different systems), rather than just the instance corresponding to the ID. With incidents as a (sub)category, potential levels of specificity could include:
  • Instance: incidents associated with a particular instance
  • Instances satisfying certain properties: incidents associated with instances whose system prompts contain harmful instructions
  • User: incidents associated with a particular user
  • A particular (type of) party: incidents involving a particular (type of) party and a system (e.g., all incidents involving hospitals and the 9-April-2024 version of GPT-4)
  • System: incidents associated with a system (e.g., the 9-April-2024 version of GPT-4)
  • Systems: incidents associated with multiple versions of a system (e.g., all versions of GPT-4)
An ID could contain (or link to) the same category of attribute at varying levels of specificity. Analogously, a tail number on an individual aircraft could be associated with maintenance records for systems, sub-systems, sub-sub-systems, etc of that aircraft.
3.2 Accessibility
The accessibility of an ID refers to who can view it and the conditions under which they can do so.
Parties that can view the ID: We distinguish between primary parties and secondary parties. Primary parties receive an ID when interacting with the corresponding instance. Secondary parties receive the ID through other means, such as directly from a primary party, or from records. Some examples include:
  • Primary: service providers; the user of the instance; other instances that interact with the instance in question [^6] Similar to how a TLS handshake allows mutual verification before establishing a communication channel, AI systems could potentially use IDs to establish trust before interacting with each other.
  • Secondary: auditors; regulators
Making ID accessible to primary parties corresponds to attaching IDs to certain outputs from the instance. Some primary parties and how IDs may be attached include:
  • Service providers: ID is included in a JSON payload
  • Users: ID is accessible through a mouseover icon in a chat interface
  • Other instances that interact with the instance in question: ID is sent in any communications with other instances
From the perspective of an actor implementing IDs (e.g., deployers), ID disclosure to secondary parties could be unintentional. For example, a data breach could render an ID accessible to the general public. As another example, a legal investigation could render IDs accessible to a government.
Selective disclosure: Some attributes may only be appropriate for certain parties. Other attributes may contain both important and sensitive information. Examples:
  • If user identification (e.g., from a know-your-customer process (Egan and Heim 2023)) is included in an ID, it may only be appropriate for government authorities to access this identification for e.g. regulatory purposes
  • A system prompt could contain both sensitive user information and information relevant for interacting parties (e.g., does the system prompt contain a jailbreak?)
Actors implementing IDs could selectively hide information from different primary parties. Those same actors could also process attributes in a privacy-preserving way (Trask et al. 2023; Sporny et al. 2024) so as to reveal only information that is relevant to the party interacting with an instance. Even so, preventing transmission of information to other parties may be difficult. For example, once an actor obtains user identification, it could—intentionally or not—transfer that information to other actors.
Persistence: IDs could be accessible for varying amounts of time after an instance has ceased to operate. Practical considerations, such as storage capacity, could impose limitations. The appropriate duration could also depend upon the application domain. Analogously, financial institutions have obligations to maintain records for set periods of time (noa 2022).
3.3 Verifiability
Suppose an author creates and sends an ID to a party that is interacting, or seeking to interact, with an instance. For the party to trust that the ID from the author provides accurate information about the instance, the following are necessary (but not sufficient; see further discussion below).
  1. (1)
    The party’s received ID is the same as the ID that the author sent
  2. (2)
    The claimed author of the received ID is indeed the author of the received ID
  3. (3)
    The received ID corresponds to the instance in question
  4. (4)
    The party trusts the author
The first three criteria lead to the following threat models:
  • Tampering: An attacker modifies the ID while it is in transit from the author to the party interacting with the instance.
  • ID spoofing: An attacker creates another ID, sends it to the party, and claims it originated from the author.
  • Instance spoofing: An attacker uses the author’s ID for their own instances.
We consider IDs (i.e., the ID system) to be verifiable if they defend against these threats. To defend against tampering, it should be easy to check if an attacker has modified the ID. To combat ID spoofing, it should be computationally difficult for the attacker to spoof authorship. For instance spoofing, it should be easy to check whether an ID indeed corresponds to a particular instance. In Sections 5 and 6, we analyze potential implementations that attempt to counteract these threats. See also (Sporny et al. 2022, 2024; Microsoft 2023), which are increasingly popular standards to implement such ID requirements in a domain-agnostic way.
Given space limitations and the complexity of the involved social questions, we do not focus on step (4) in this work. Similarly, we also scope out issues relating to the accuracy of the attributes in an ID. Accuracy is not the same as verifiability, since even if (1)-(4) are satisfied, the author could have made mistakes in creating and linking to the information.
4 Demand for ID Use
IDs seem useful for actors engaging with an AI system, but why would a user or deployer of that AI system choose to present an ID? In short, governments, service providers, and parties interacting with AI systems all have interests in ID use, as well as means to incentivize (or mandate) it.
4.1 Governments
Interests in ID use: Governments may wish to disincentivize and investigate harms caused through AI systems. In the case of a financial scam for instance, the IDs of AI systems involved in any transactions could aid investigation. Similarly, governments have an interest in preventing fraud, and therefore already mandate that certain financial transactions involve ID checks.
Mandating ID use: Governments could mandate that certain, highly consequential service providers obtain IDs from any AI systems with which they interact. For example, financial institutions could request IDs when AI systems make large financial transactions.
4.2 Service Providers
Interests in ID use: IDs could be advantageous for service providers. First, IDs could help to reduce service abuse. Plugins and services—as well as the unique vulnerabilities of AI systems, such as prompt injection attacks (Zhan et al. 2024; Greshake et al. 2023)—could allow attackers to harm users or the service provider (Stumpp 2023). As a countermeasure, service providers could use IDs to block abusive instances. Since attackers could avoid block lists by creating instances with new IDs, service providers could limit the privileges of instances they have not interacted with before, and provide preferential services to certain, trusted IDs (e.g., instances with a history of good behavior; see more below). IDs could also aid own incident investigation processes. Second, service providers could use IDs to adapt services to the usage patterns of particular instances, which could help to retain service users and reduce costs.
Incentivizing ID use: Service providers could develop, or encourage the development of, plugins that require IDs. AI systems can already perform a variety of useful tasks through plugins, including web searches, email communication, and stock trading (OpenAI 2023a; Richards 2023; Wu et al. 2023; Anthropic 2024b). In comparison to an AI system mimicking a human’s interaction with a computer (e.g., interacting with individual web page elements), plugins may be more reliable and performant for users. AI systems continue to have difficulty with the former (Furuta et al. 2024; Tao et al. 2023; Gur et al. 2024, 2023; Xie et al. 2024), but plugins already enable useful tasks. As well, a lack of constraints with such mimicry—e.g., accidental file deletion from mouse/keyboard operation—could result in more severe problems than with plugins, which constrain the actions available to the AI system.
Yet, any actor could write a plugin for a service that does not require IDs. A potential response could be for service providers to restrict their services in the absence of an ID. Restrictions—such as rate limits, throttling, or quotas—already reduce abuse of many existing services (Cloudflare 2024). The design of restrictions could still allow humans to interact with services while limiting AI systems without IDs to human speeds, so as to reduce the impact of potential abuse. On the other hand, service providers could provide additional privileges to AI systems with IDs. This selectivity could incentivize ID use, since systems that presented IDs would have performance advantages.
Other actors could bypass plugins in any case by having an AI system mimic human interaction with a computer. A potential countermeasure could be CAPTCHAs-like methods to distinguish between human and non-human actions. Although software systems have become increasingly capable of bypassing CAPTCHAs, [^7] Task-based CAPTCHAs are no longer as common as CAPTCHAs based on behavior patterns, such as how a mouse is moved across the screen (Shet 2014). Yet, AI systems may one day be able to imitate human behaviour patterns as well (Adept 2022). they could be an interim option in anticipation of more robust proof-of-personhood protocols (Buterin 2023). Entities that failed CAPTCHAs could be subject to ID requirements, or else face service restrictions.
4.3 Other Parties that Interact with AI Systems
Interests in ID use: Some parties, such as businesses, may only wish to interact with certain, trusted AI systems when performing transactions. IDs could provide the information to identify such systems. On the other hand, it is unclear if individual consumers would use IDs. Yet, future software—including users’ personal AI assistants (Gabriel et al. 2024)—could inspect IDs and reject potentially unsafe interactions (e.g., with jailbroken systems), similar to how browsers warn users before visiting websites without valid HTTPS certificates.
Incentivizing ID use: Large corporations could exert strong pressure for ID usage, through reducing or avoiding engagement with AI systems without IDs. Future software that inspects IDs could also exert similar pressure. Analogously, websites that do not use HTTPS are disadvantaged since popular web browsers warn users when they are making an insecure connection.
4.4 Where to Require IDs
While IDs could be useful as described above, universal ID requirements might not be warranted. As we will discuss in Section 7, ID implementation imposes burdens and risks, and there remains uncertainty about the broader impacts of ID usage. As such, if governments or service providers decide to incentivize or require ID use, they should target where IDs could be most useful, such as high-stakes settings. We provide some examples of high-stakes tasks which could justify IDs.
Making large financial transactions: Some service providers could allow AI systems to perform financial transactions, such as making purchases, trading stocks (khan 2024), or otherwise transferring funds. Such service providers could require IDs for transactions involving significant sums. IDs could help to reduce the risk of a variety of negative outcomes, by enabling service providers to filter out potentially untrustworthy AI systems (e.g., those that may likely malfunction and make mistaken transactions) and investigate potential incidents.
Healthcare interactions: For better or for worse, some users are already turning to chatbots for health advice (Lucas 2024). IDs could help users make informed choices about such interactions. If AI systems mediate increasingly many medical interactions (Singhal et al. 2023), IDs could become more useful for establishing trust. In response to current trends, a variety of organizations—such as medical device manufacturers and hospitals—could provide plugins to connect AI systems to users’ healthcare data. Patients could have personal assistants that communicate with the AI systems of healthcare providers, so as to facilitate diagnosis and treatment.
5 ID Implementation for Centralized Deployments
We assess a basic ID system that deployers of AI systems could implement. Our discussion in this section will be agnostic to the applications where IDs are used (see Section 4.4 for discussion of where ID use may be warranted). In Section 6, we explore how IDs could work in more decentralized deployments.
We assume a straightforward choice for the identifier, system_identifier:instance_identifier , where system_identifier is the same for all instances of a system [^8] If the underlying model changes, the system ID should also change. E.g., ChatGPT with a GPT-4 backend is different from ChatGPT with a GPT-3.5 backend. and instance_identifier is unique to a particular instance. It could be useful and straightforward to append other details to the identifier, such as the exact time at which the instance was created. However, we proceed with this choice for simplicity.
5.1 Assigning IDs
Although IDs are instance-specific, it may not be feasible for deployers to detect the formation of all new instances. For example, a user could use deployer-run instances $A$ and $B$ as backends for a user-run system $C$, which integrates the outputs of $A$ and $B$ and would evade detection by deployers. We outline some scenarios where deployers could detect new instances and analyze considerations for assigning new IDs.
New instances: Deployers can easily generate new identifiers whenever users create new instances (e.g., a “new chat” button).
Reloaded instances: Suppose that a deployer saves instances’ states (i.e., the history of interaction) and provides separate functionality for users to reload them. We tentatively suggest that deployers retain the ID of the original instance for the reloaded instance, since some attributes (e.g., prior incidents, context) of the original instance could be relevant for understanding the behaviour of the reloaded instance. Alternatively, if a deployer does decide to assign a new identifier to the reloaded instance, a link to the ID of the original instance would maintain accessibility to the relevant attributes.
Output regeneration: Some deployers provide the ability for a user to regenerate an output in response to previous input. Although not yet implemented, deployers could additionally allow users to create multiple branches (see Appendix A for further discussion of branches) of interactions, based off of regenerated outputs. More concretely, a user could interact with their instance up until the output $a_{t}$ at time $t$, regenerate to obtain an output $b_{t}$, and continue interactions with the respective branches to obtain outputs $(a_{t+1},b_{t+1})$, $(a_{t+2},b_{t+2})$, etc. Since the two branches have no necessary causal impact on each other (unless they interact with each other or modify a shared object in the world), they should have two separate IDs, with links to the ancestor ID.
Composite systems: Software frameworks could give users the ability to create new instances (a composite) out of a collection of existing instances (Wu et al. 2023). If deployers run such software frameworks for users, they could detect the formation of composites. Since the constituent instances can continue running alongside the composite, composites should have new IDs, but could have their IDs linked to their constituents’.
Fine-tuned systems: Some deployers provide fine-tuning functionality to users. Since fine-tuning can change the behaviour of a system, a fine-tuned system $A^{\prime}$ should have different system_identifier than the original system $A$, and link to $A$’s system_identifier . Although it may be useful to include some documentation about the fine-tuning data in an ID, user privacy expectations could complicate such inclusion. Some deployers monitor fine-tuning data only for abuse detection and legal compliance (ChrisHMSFT et al. 2023; OpenAI 2024a); inclusion of fine-tuning data in an ID likely requires stronger justification.
5.2 Attributes
We tentatively suggest the following attributes for a deployer to include (or link to) in an ID:
  • A system card
  • A database of incidents associated with the system
  • The IDs of ancestor and descendant instances
If instances of future systems persist longer in the world, it may be appropriate to link to a (sub)database of incidents associated with the instance. Implementing links to ancestor and descendant instances, when feasible in the situations discussed in Section 5.1, could be a simple way to test the utility of instance-level identification.
There are likely ways for users to avoid ancestor or descendant links. For example, a user could manually copy the inputs of a previous instance to a new instance, avoiding a deployer-provided reloading function. More reliable identification of ancestors and descendants could involve more invasive measures, which may only be appropriate in high-stakes domains. For example, a deployer could attempt to analyze the inputs of different instances from the same user for similarities.
5.3 Accessibility
Users: The ID could be accessible through a user interface and sent along with any API requests from a user. If a deployer serves an API to a customer, who subsequently serves an API based on the deployer’s system to another user, the deployer would likely have to work with the customer to ensure that the end user receives an ID (see Appendix B for further discussion).
To help to ensure that secondary parties maintain access to the ID, the deployer could include IDs in watermarks and metadata of media outputs (i.e., text, images, video). Yet, it is unclear how robust such methods may be to user removal (Zhang et al. 2023; OpenAI 2024c).
Services: When the AI queries a service, the ID should be sent along with the request. Since requests are text-based (e.g., JSON), including the identifier and (links to) the attributes would be straightforward.
5.4 Verifiability
To ensure verifiability, we explore an approach similar to that of digital certificates:
  1. (1)
    When an instance is created (e.g., when a new chat session starts), the deployer generates an ID for the instance.
  2. (2)
    Whenever the instance generates an output, the deployer attaches a digital signature for the combined ID and the output (e.g., via hashing the output).
  3. (3)
    The deployer sends the ID, signature, and output to the party interacting with the instance.
We emphasize that an ID would be associated to outputs via digital signatures. Hence, each output would have a different signature. Even though the information contained an ID might not change between different outputs, the signature verifying that the ID corresponds to an instance’s output would. This step is crucial to prevent instance spoofing. Otherwise, any actor could copy an ID, generate an output from an arbitrary instance, and present the copied ID as that of the arbitrary instance. [^9] Note that the signatures on HTTPS certificates do not need to be based on website content. Even if an attacker copied the details of an HTTPS certificate, they would not have the private key necessary for establishing a secure connection with a user. A digital signature also makes tampering and ID spoofing computationally difficult.
Deployers could sign outputs at differing granularities. For example, given a text output, should a deployer generate a signature for each token of the output? As a general principle, we suggest that the to-be-signed output should at least contain a functional unit for the party receiving the ID. For example, suppose a service provider receives requests in a JSON string. A signature of the ID and closing brace (“}”) of the JSON string would allow the attacker to replace the payload of the JSON string while maintaining the validity of the signature. On the other hand, a signature of an ID and the entire JSON payload would avoid this failure mode. We explore some additional examples of applying this principle:
  • Users: Deployers could sign the outputs generated since the last user input.
  • Other parties interacting with the instance: The deployer could sign the output which it expects the party to see. For example, if deployers provide a way for different instances to communicate with each other, the deployer should sign the sent messages.
However, requiring signature of the output comes with a crucial caveat. If any actor were to change or further process the output of an instance, the digital signature would not be valid for the modified output. Much flexibility could be lost under this implementation, if IDs were to be required in many settings.
Part of preventing ID spoofing involves guaranteeing that the digital signature indeed belongs to the deployer. Existing public key infrastructure for doing so involves a trusted third party known as a certificate authority (CA), who certifies that public keys (which allow third parties to verify digital signatures) belong to stated parties (the deployers in this case). The existing system of CAs could support the authenticity of the deployer’s signature. Alternatively, a collection of civil society and industry organizations could act as an initial CA, and delegate authority to certain deployers after having verified their identities. Future work should explore these options further, as well as decentralized alternatives to CAs that avoid concentrated points of failure (Sporny et al. 2022, 2024).
Figure 2: This figure illustrates the process in Section 5.4 of creating IDs that are verifiable.
6 ID Implementation for Decentralized Deployments
In some settings where ID use may be warranted (e.g., high-stakes settings; see Section 4.4), users themselves may run AI systems. We assess an implementation of IDs without deployers as an intermediary, which would require more involvement from users, CAs, and supporting software. As in Section 5, we adopt system_identifier:instance_identifier as the identifier.
6.1 A Basic Design
This design roughly follows that in Section 5, except that we will have AI frameworks—software tools such as the transformers library (Wolf et al. 2020), AutoGen (Wu et al. 2023), and AutoGPT (Richards 2023) that facilitate running AI systems—facilitate ID implementation. Similarly, automatic certificate issuance from Let’s Encrypt played an important role in the widespread adoption of HTTPS (Encrypt 2024). AI frameworks will require some additional functionality that does not yet exist, as we explain below.
  1. (1)
    The user registers for an account with a CA.
  2. (2)
    The AI framework lets a user log into their CA account.
  3. (3)
    Whenever a user starts an instance with an AI framework, the framework requests an identifier from the CA.
  4. (4)
    For each unit of output from the instance, the AI framework encrypts the output and sends it to the CA.
  5. (5)
    The CA produces a digital signature for the combined identifier and hashed output.
  6. (6)
    The AI framework makes the ID and signature accessible with the output.
As in Section 5.4, signature of the output would prevent attackers from attaching a given ID to arbitrary instances. To ensure user privacy, the CA would only sign a hashed output. Considerations for which outputs to sign (e.g., signing the entire JSON string of a service request) from Section 5.4 apply here as well.
As a potential way to ensure that CAs do not sign arbitrary outputs from users (e.g., those not even generated by an AI system), CAs could stipulate the use of approved AI frameworks. Techniques such as remote attestation and code signing could be useful for such requirements. [^10] Such techniques would ideally only apply to the loading and use of relevant library packages, rather than all parts of a user’s code base. Yet, only permitting the use of approved frameworks could have negative consequences, especially if IDs were required in many settings. Users would not be able to modify frameworks as they see fit. Bottlenecks in the approval process could also mean that users are pressured to use outdated frameworks, or frameworks that do not provide desired functionality.
Figure 3: This figure illustrates a potential process for ensuring verifiability of identifiers in decentralized deployments. See further details in Section 6.1.
Information to include: Some AI frameworks, such as HuggingFace, integrate with repositories or marketplaces where users can download different AI systems for use. Frameworks could link to the documentation that is often included in a system’s repository. However, unlike in Section 5.2, there is no guarantee that a repository would contain a system card or an incident database. If the AI framework provides a dedicated function for an ancestor instance to create descendant instances, the AI framework could establish ID links between the ancestor and the descendant. Otherwise, such links seem difficult to establish.
Accessibility: An AI framework has a similar role to that of a deployer in Section 5.3. The framework should ensure that any outputs of the system, including audiovisual outputs and queries to services, should contain the ID, whether embedded as a watermark or as metadata.
Verifiability: Verifiability would apply to the identifier, but not necessarily to other information attached to the ID. On HuggingFace for example, any anonymous user can upload a system with arbitrary documentation. Even so, such lack of verifiability may encourage parties to interact with instances of well-known systems, rather than instances of systems uploaded by anonymous users. This preference could incentivize users to use well-known systems and weaken the need for additional verifiability mechanisms.
6.2 Assigning IDs
In theory, AI frameworks could detect instance creation similarly to how deployers could in Section 5.1, and request new identifiers from CAs as appropriate. The main practical limitation is the flexibility that users have—by virtue of writing and running their own code—to create instances in arbitrary ways. It may not be feasible for an AI framework’s developer to account for all potential instance creation methods in advance, and decide how to assign IDs accordingly.
7 Limitations
We discuss limitations of IDs and of our analysis.
7.1 Failure to Inform
Parties could ignore IDs, just as internet users sometimes ignore indicators of untrustworthiness (Norris and Brookes 2021). Software—or future AI assistants (Gabriel et al. 2024)—could make informed decisions for users in some circumstances, just as browsers reject websites with invalid HTTPS certificates. However, it remains unclear what information would justify rejecting an interaction with an AI system. Other parties could misunderstand or misinterpret an ID’s attributes. For example, it can difficult to assess the external validity of evaluation results (Fourrier et al. 2023; Weidinger et al. 2023).
Moreover, ID attributes could be inaccurate. For example, deployers could unintentionally link to an incorrect version of system documentation, or may simply fail to update the documentation when a system is updated. OpenAI has occasionally released new versions of GPT-4, [^11] See https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4. but has not updated the GPT-4 system card (OpenAI 2023b) since the initial release, as of April 2024. Regular audits (Raji et al. 2020; Sharkey et al. 2023) could verify the accuracy of attributes.
7.2 Bypassing IDs
As with other types of real-world identity systems, IDs for AI systems cannot wholly prevent undesirable behavior. IDs could be lent, or the benefits obtained from ID use—such as purchase of a good—could be transferred. Users may hesitate to lend IDs since the borrower could inflict lasting reputational damage on the ID. Yet, similar to the purchase of alcohol for minors, it seems difficult to track or prevent transfer of goods. Future work could assess how likely and how damaging such transfer could be.
7.3 Limited Adoption
In decentralized deployments, the absence of supporting AI frameworks could hinder users from implementing IDs. Without such software, providers may be hesitant about requiring IDs if they wish to encourage use of their services. Consequently, there may not be enough demand to spur the development of supporting software. One way to resolve these problems is to focus on soft ID incentives with well-resourced actors. For example, service providers could experiment with ID incentives for companies, who could pilot initial implementations of supporting AI frameworks.
In the implementations we discussed, verifiability requires a signature on both the ID and the output. Modification of the output would hence invalidate the ID. Users could be hesitant to adopt IDs given this lack of flexibility. As well, many generated IDs could become invalidated simply because users unintentionally modify the outputs (e.g., a user accidentally removes some punctuation from generated text). ID requirements could be less onerous when outputs are unlikely to be modified, such as requests from an AI system to a service. Yet, absent technical implementations that can ensure verifiability while permitting modification, this limitation suggests that any initial ID requirements should be minimal.
7.4 Privacy and Security Risks
Although an identifier by itself would not reveal anything about its user, the user’s privacy could still be compromised in a number of ways. First, information about the instance could reveal user details. For example, timestamps and details of an instance’s activities—as might be contained in incident reports—could help a third-party to identify users. The existence of IDs may engender pressure to attach additional information, which could facilitate further user identification. For example, service abuse could potentially be correlated with certain activity patterns, which could motivate corresponding data collection. Second, if the deployer maintains an internal database linking a system’s ID to the corresponding user account, other parties could obtain and misuse the information. Security vulnerabilities could allow attackers to obtain the database. Overreaching governments could force deployers to reveal the users of particular systems.
In the context of the technical implementations we discussed, the trustworthiness and security of the CAs limits the verifiability of IDs. Apart from external threats, high-stakes situations could create incentives for collusion between authorities and key owners. With a compromised CA, a malicious actor could issue IDs for any AI system. For example, an attacker could assume the ID of a trusted AI system so as to interact with certain parties or obtain higher privileges with a service.
Decentralized alternatives to digital signature management could help to avoid these problems. Specifically, smart contracts deployed on distributed ledgers could potentially issue digital signatures without the need for a centralized authority (Omote 2023). The smart contracts would be publicly verifiable and are immutable as long as the integrity and availability of the distributed ledger as a whole, in contrast to a single trusted authority, is ensured. Yet, widespread adoption of such approaches is currently limited. Challenges include a lack of usable and tested software products, scalability and compute demands, and technical issues inherent in real-world implementations of distributed ledgers.
7.5 Broader Societal Consequences
Since IDs could enable anybody to identify and make decisions about particular instances, their use could have broad societal consequences.
IDs with additional user information could enable potentially harmful ranking of users. Actors implementing IDs may wish to include such information if it is correlated with undesirable outcomes, such as fraud. Yet, the existence of confounders could lead to instances of some users being unfairly denied access to services, similar to how toxicity classifiers can biased against African American English (Sap et al. 2019). Even if the relationship between information and undesirable outcome was causal, taking such information into account could still be unjustly discriminatory.
IDs could also enable influence over particular instances and their users. Even if an ID does not contain information about an instance’s interactions, external actors could record particular IDs have been involved in interactions, just as cookies can track user activity across websites. Instances that are active for long periods of time could reveal much information about a user’s preferences. Actors interacting with instances could try to influence their behaviour through means such as prompting. Advertisers and businesses could try to get instances to purchase or recommend their products or services for their corresponding users. Political actors could try to get instances to influence the voting behaviour of their users. Governments could attempt to enforce rules upon particular instances. Such influence could be strong, especially if personal AI assistants become much more widespread and central to daily interactions (Gabriel et al. 2024). The specific impacts of this influence deserve further study.
Finally, IDs could create a separate digital channel for interactions between trusted AI systems. The vast majority of digital interactions could in the future be between AI systems, especially if future AI assistants handle most of a user’s activities (Gabriel et al. 2024). If IDs enable reliable interactions, users may prefer their AI systems to interact only with AI systems that present IDs. Much digital activity could take place between networks of ID-bearing AI instances that trust (and can verify) each other. Entities that trust each other may be able to engage in more productive interactions. As a corollary, instances without IDs may become severely disadvantaged, relative to instances within ID-bearing networks. To avoid such disadvantage, users may have no choice but to submit to the potential negative consequences of IDs we have discussed heretofore in this section. On balance, the consequences of a separate digital channel for AI systems remain unclear.
8 Related Work
Besides real-world examples of IDs and work on documenting AI system that we highlighted in Section 1, we discuss two additional lines of related work.
A growing line of work is concerned with building digital infrastructure to structure how AI systems, especially agents (Chan et al. 2023; Shavit et al. 2023; Gabriel et al. 2024), behave and interact. Such infrastructure can have a variety of goals. Patil et al. (2024) build a runtime for LLM agents to enable human validation and reversal of actions. Marro (2024) sketches a protocol for communication between LLMs; the protocol relies on documents to specify structured rules, with natural language as a fallback. (Sun et al. 2023) argue that decentralized commitment devices will be necessary to allow agents to coordinate with each other. IDs are a form of infrastructure that enable informed decision-making about AI interactions.
The technical implementations we assess take inspiration from several Internet and security protocols. HTTPS (Fielding, Nottingham, and Reschke 2022) uses digital certificates (Cooper et al. 2008) to allow users to verify the identity of an accessed website. If the certificate is valid, a user’s browser uses the information contained in the certificate to establish a secure connection (Rescorla 2018). Websites must obtain such certificates from CAs (Cooper et al. 2008), who verify the identity of the website owner and issue the certificate upon successful identification. Identity verification depends upon public key cryptography (Barnes et al. 2019), which allows a party to prove ownership of a given, public identifier, otherwise known as a public key. Let’s Encrypt is a non-profit CA which automates the process of issuing certificates (Encrypt 2024). HTTPS everywhere was a browser plug-in that forced the usage of the HTTPS version of websites whenever it was available; this functionality is now default in modern web browsers (Hancock 2021). To design further digital infrastructure for AI systems, other internet and security protocols may be a fruitful source of inspiration.
9 Conclusion
To inform crucial decisions about AI interactions, we proposed a framework for creating IDs for AI systems. IDs can vary in their attributes, how accessible they are to various parties, and to what extent they are verifiable. There could be significant demand for IDs from governments, service providers, and users, particularly when systems engage in high-stakes tasks. These actors also have means to incentivize ID usage. ID implementation seems feasible for deployers, but implementation in decentralized settings will require further study.
More research is required to understand and address the potential risks of IDs. First, security and privacy risks should be addressed, such as the potential for certificate authorities to become compromised. Second, it is unclear under what circumstances IDs should be accompanied by user identification. Third, the broader societal consequences of introducing IDs require further study.
Limited experimentation with IDs seems justified given their potential benefits and risks. IDs seem most warranted in high-stakes settings, where certain actors could experiment with incentives for ID use. For instance, service providers that provide plugins in potentially high-stakes settings (e.g., plug-ins that allow AI systems to make financial transactions) could experiment with rate-limiting instances without IDs, while still allowing such instances to access services. Deployers could begin to implement the technical infrastructure for IDs.
IDs have facilitated essential functions across a variety of domains. Incident investigation, allocation of liability, and establishment of trust would be more difficult without IDs for products, organizations, and software systems. As AI systems become increasingly prevalent, deficiencies in such functions could result in more severe consequences for society. If implemented well, IDs for AI systems could enable mechanisms to avert such consequences and navigate this emerging world.
Appendix
Appendix A A More Detailed Definition of an Instance
A system is a model (e.g., a set of parameters), along with software used to run the model and provide other user functions. For example, ChatGPT with GPT-4 as a backend comprises both the weights of (a particular version of) GPT-4 and the software used to facilitate chat interactions (as opposed to an “autocomplete” function, as in the OpenAI playground).
To understand the difference between a system and an instance, consider the information that would be useful to attach to a system-specific ID, so as to inform decisions about interacting with the system. Such information could not depend (too much) on a system’s inputs, since different inputs might lead to different behavior. In other words, the information should ideally be valid regardless of a user’s interaction history. Represented visually, the information would have to be valid for both top and bottom flow charts in Figure 4.
Figure 4: This figure depicts two users that use the same system (e.g., both use ChatGPT with the GPT-4 backend). System-specific information attached to an ID should be useful to both users. By loaded, we mean that the system is ready to accept inputs for the first time.
An instance is an abstraction that corresponds to a (initial) user (which could be a human, a group of humans, a software system, etc) and an interaction history. In Figure 5, we provide a visual depiction of two instances that interact with each other.
We define instances in this way so that instance-specific IDs can take into account information that is causally relevant to a given interaction. [^12] There may be information that is not causally relevant, but which may still help to predict the instance’s behavior, such as model evaluations or the behaviour for other instances. It remains unclear how much of this information to attach to a given instance’s ID. For example, in the top half of Figure 5, information about instances $A$’s earlier interactions (such as malfunctions) may be useful when instance $B$ interacts with instance $A$. Actions taken by another instance are not, by default, causally relevant to the behaviour of another instance. Yet, instances can affect each other through direct interaction or on changing shared states of the world (e.g., using the same bank account).
Figure 5: An instance is an abstraction that corresponds to a creation event, where a system is loaded for a user, and an interaction history. Instance-specific IDs could help to inform interaction decisions. For example, information about instances $A$’s earlier interactions (such as malfunctions) may be useful when instance $B$ interacts with instance $A$.
Since AI systems can be copied and combined, there are some additional edge cases for how to define instances. In a branch (see Figure 6(a)), Two instances share a past context. For example, a regeneration of a response (possible in ChatGPT, for one) creates two branches. Since branches can behave independently, we suggest treating branches as separate instances. In a merge (see Figure 6(b)), two instances come together to form a new system. For example, suppose software $S$ accepts inputs from separate users, engages the instances of two users in a debate based on the inputs, and finally outputs a result to both users. Users $A$ and $B$ could use $S$ to create a new instance $AB$, based on their separate instances. As with branching, we treat $AB$ as a separate instance because the users could continue running their original instances in tandem with $AB$.
(a) In this example of a branch, the inputs of instance $A$ are copied for user $C$ to create a new instance $C$.
(b) In this example of a merge, the instances $A$ and $B$ are merged into a new instance $AB$, which has access to (potentially a subset of) the inputs of $A$ and $B$.
Figure 6: We illustrate how we define instances in the event that they are copied or combined.
If an instance is the result of a branch or a merge, it may be useful for the instance’s ID to contain information about ancestor instances. In Figure 7(a), since the context of instance $A$ is copied into instance $C$, any incidents associated with $A$ may be helpful for user $B$’s interaction decisions. Similarly, in Figure 7(b), information about instances $A$ and $B$ may be relevant for user $C$’s decision about whether or not to trust the outputs of instance $AB$. Nonetheless, information from ancestor instances could be excluded for a variety of reasons, such as user privacy. Furthermore, it could be possible for information from descendant instances, or instances in another branch, to be relevant for the behavior of a given instance. For example, incidents with user $B$ from time $t+1$ may be informative for user $A$ that interacted with an instance at time $t$. We leave further analysis of how to delimit the sources of ID information to future work. [^13] Links to ancestor IDs make other potential definitions of an instance equivalent, from the perspective of the functions of an ID. For example, in Figure 6(a) we could have defined instance $C$ as encompassing interaction history up to instance $A$. This definition would not change the causally relevant information with respect to instance $C$ in Figure 7(a).
(a) In the event of a branch, a user $B$ that interacts with instance $C$ may find information about instance $A$ (e.g., instance $A$ was involved in an incident) to be useful.
(b) In the event of a merge, a user $C$ that interacts with instance $AB$ may find information (e.g., instance $B$ is not robust) about either instances $A$ or $B$ to be helpful.
Figure 7: We visually represent the information that an ID for an instance could take into account.
As we have defined it, an instance is causally independent of other instances unless it interacts with them, such as by processing their outputs. This causal independence is crucial to IDs, as we discussed above. Yet, the abstraction of an instance could be leaky depending on system implementation details. For example, if a single GPU handles the operation of two separate instances, implementation errors or optimization strategies (e.g., approximate matrix multiplication) could potentially result in computational interference between the instances.
Appendix B Maintaining the Accessibility of IDs
Suppose a primary party $P$ receives an output and ID from an AI system. If it is desirable for secondary parties to see the ID, $P$ may have additional responsibilities. Some potential situations follow.
Chains of deployers: There may be multiple deployers involved in the operation of an AI system. For example, deployer $A$ could provide an API for system $X$ to customer $B$, who creates system $Y$ and serves $Y$ to user $C$. To ensure that $C$ can access the $X$’s ID, $B$ should link to the ID provided by $A$, or should include the information in the ID of $Y$.
Service providers: If a service provider receives an ID, it should ensure that those who observe or are affected by AI system’s actions can also see the ID. For instance, a bank that facilitates a financial transaction from an AI system should make the latter’s ID known to the other party of the transaction.
The provider may have to work with other parties to maintain visibility of the ID. As an example, consider a service that allows an AI system to post on a social media platform. The provider may have written the software that allows the AI system to interact with the platform’s API, but the social media company develops the API and manages the platform. While the provider could include the AI system’s ID as text within the post, doing so may be infeasible because of text limits or would otherwise be obstructive to the user experience. Rather, the provider should work with the social media company to include the ID as metadata, readily accessible to the user through a visible icon.
Users: Users could receive outputs with attached IDs, whether as a watermark or metadata. To counter inadvertent removal of IDs, deployers could add labels to inform users. For example, Facebook and Instagram automatically add labels indicating AI origin when a user shares a photo generated with the Meta AI feature (Bickert 2024). To counter intentional removal of IDs, better watermarking techniques may be required. Yet, it remains unclear how effective watermarking may ultimately be (Zhang et al. 2023).